Question 10 Not yet answered Points out of 4.00 P Flag question Do bonds reduce the overall risk of an investment portfolio? Let x be a random variable representing annual percent return for the Vanguard Total Stock Index (all stocks). Let y be a random variable representing annual return for the Vanguard Balanced Index 60% stock and 40% bond For the past several years, assume the following data. Compute the coefficient of variation for each fund. Round your answers to the nearest tenth. 13 38 23 35 25 26 -13 -13 16 26 16 16 -2 -3 Select one a. for x-values: 194.1%, and fory-values: 236.8% b. for x-values: 108.6%, and fory-values: 236.8% c. for x-values: ¡08.6%, and fory-values: 132.4% d. for x-values: 132.4.6 and fory-values: 194.1% e. for x-values: 194.1%, and fory-values: 132.4% Previous page Next page
Answer #1

Solution:

Coefficient of variation can be computed as

Coefficient of variation =(SD/mean)×100%

Mean (X) = 11.8

SD(X)= 20.7353

Mean(Y)= 9.1

SD(Y)= 12.05

So CV(X) = (20.7354/11.8)×100%= 174.04%

CV(Y)= (12.05/9.1)×100%=132.4%